.A coal power plant consumes 100000 kg of coal per hour and produces 500 MW of power. If the heat combustion of coal is 30 MJ/kg, what is the efficiency of the power plant?


1
Expert's answer
2021-10-10T16:10:13-0400

Efficiency in terms of power

"e=\frac{P_0}{P_i}"

The power input by the coal is

"P_i =AH \\

P_i = (10^5 \;kg/hr)(30\;Mj/kg) \\

= 30 \times 10^5 \;mJ/hr \\

P_i = [30 \times 10^5 \;mJ/hr][\frac{1 \;hr}{3600 \;s}] \\

= 833.3 \;Mj/s \\

= 833.3 \;MW"

The efficiency of power plant is

"e =\frac{P_0}{P_i} \\

= \frac{500 \;MW}{833.3 \;MW} \\

= 0.6"